OpenStudy (anonymous):

6x = 7 – 4y (1 point)

OpenStudy (angela210793):

What to do with this? O.o

OpenStudy (anonymous):

what to do??

OpenStudy (anonymous):

6x=7-4y 6x-7=-4y -6/4x+7/4=y y=-3/2x+7/4 Yes, this is a function.

OpenStudy (anonymous):

i hope thats wat u wanted! :)